OpenStudy (anonymous):

Find the dimensions of largest rectangle w/ maximum area inscribed in a semicircle with radius r. What equations am I gonna use in here? Help!

OpenStudy (anonymous):

The equation of the semicircle is \[x^2+y^2=r^2\] The area will be \[A=2xy\] So \[y=\sqrt{r^2-x^2}\] and \[A(x)=2x\sqrt{r^2-x^2}\]

OpenStudy (noelgreco):

It's now time to use differential calculus so that you can maximize A. What is dA/dx?

OpenStudy (anonymous):

\[\frac{dA}{dx}=2\sqrt{r^2-x^2}-\frac{2x^2}{\sqrt{r^2-x^2}}=\frac{2r^2-4x^2}{\sqrt{r^2-x^2}}\] \[\frac{2r^2-4x^2}{\sqrt{r^2-x^2}}=0\] \[x=\frac{r}{\sqrt{2}}\]